Exercise 9-20 Sunland Company purchased a new machine on October 1, 2017, at a cost of $86,600. The company estimated that the machine has a salvage value of $8,400. The machine is expected to be used for 65,700 working hours during its 8-year life. Compute depreciation using the following methods in the year indicated. Declinng-balance using double the straight-line rate for 2017 and 2018. (Round answers to 0 decimal places, e.g. 125) 2017 2018 Depreciation using the Declining-balance method s Calculate the depreciation cost per hour. (Round answer to 2 decimal places, e.g. $1.25) Depreciation cost per hour Units-of-activity for 2017, assuming machine usage was 500 hours. (Round answer to 0 decimal places, e.g. 125) Depreciation using the Units-of-activity method for 2017 Click if you would like to Show Work for this question: Open Show Work
Answer #1

Answers

  • All working forms part of the answer
  • Working for Declining balance method:

A

Cost

$            86,600.00

B

Residual Value

$              8,400.00

C=A - B

Depreciable base

$            78,200.00

D

Life [in years]

8

E=C/D

Annual SLM depreciation

$              9,775.00

F=E/C

SLM Rate

12.50%

G=F x 2

DDB Rate

25.00%

  • Answer #1:
    Depreciation using Declining balance method:

Year 2017 = $ 86,600 x 25% x (3 months/12 months) = $ 5,413
Year 2018
= ($86600 – 5413) x 25% = $ 20,297

  • Answer #2:
    Depreciation cost per hour = Depreciable base / Expected working hours

= $ 78,200 / 65,700 hours

= $ 1.19 = Answer

  • Answer #3:
    Depreciation using Units of activity method for 2017 = 500 hours x $ 1.19 per hour

= $ 595 = Answer